OpenStudy (anonymous):

Two vertices of a right triangle exist at (3,2) and (6,7). Find the x-coordinate of the third vertex given that the y-coordinate of the third vertex is -8. I got that the x-coordinate is 59/3 is that right?

OpenStudy (anonymous):

@phi help!

OpenStudy (phi):

you can check by finding the slopes of the two legs, they should be negative reciprocals or find the lengths of all 3 sides and see if pythagoras theorem works.

OpenStudy (anonymous):

@phi but it says there are two solutions and i can only get x=31

OpenStudy (anonymous):

i did the negative reciprocal thing and it only gives x=31 as a solution

OpenStudy (anonymous):

i found the slope to be 5/3

OpenStudy (phi):

the slope of the line through the points (59/3, -8) and (3,2) is -3/5 the slope of the line through the points (3,2) and (6,7) is 5/3 they are negative reciprocals, so those two lines intersecting at (3,2) for a right angle

OpenStudy (anonymous):

then i did:\[-\frac{ 3 }{ 5 }=\frac{ -8-7 }{ x-6 }\] and solved for x

OpenStudy (anonymous):

but the problem says 2 solutions?

OpenStudy (phi):

we could have the right angle at the other point (6,7)

OpenStudy (anonymous):

wait but when i solve this equation i get 31 as a solution

OpenStudy (anonymous):

for the xcoordinate

OpenStudy (phi):

yes, you put the right angle at point 6,7

OpenStudy (anonymous):

How do you do it if the right angle is at (6,7)?

OpenStudy (anonymous):

oh so the two solutions are: 31 and 59/3?

OpenStudy (phi):

if you do the same thing but at (3,2) you will get the other solution

OpenStudy (anonymous):

so i did it all myself without knowing it

OpenStudy (anonymous):

so i am right?

OpenStudy (phi):

yes, (31,-8) is one of the solutions

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

I get confused sometimes

OpenStudy (phi):

here is a graph of the problem

OpenStudy (phi):

the slope between (3,2) and (6,7) is 5/3 you want the slope from (3,2) to (x,-8) to be -3/5 or \[ \frac{-8 - 2}{x-3} = \frac{-3}{5} \\ \frac{-10}{x-3}= \frac{-3}{5} \\ \frac{10}{x-3}= \frac{3}{5}\] flip both fractions and then multiply by 10 \[\frac{x-3}{10}= \frac{5}{3} \\ x-3 = \frac{50}{3}\\ x= 3+\frac{50}{3}= \frac{59}{3} \] is the other solution, (which you already found)